# Improving Robustness and Reliability in Medical Image Classification with Latent-Guided Diffusion and Nested-Ensembles

### v4

@@ III-B Training
Step 1 Training TE Blocks.  The parameters of the TE blocks are learned during the training of the ViT. As discussed previously, the input $\mathbf{x}$ is first transformed into the embedding $e_{0}^{\text{wcls}}$, which is then passed through $L$ TE blocks to produce $e_{L}^{\text{wcls}}$. Finally the class token $e_{L}^{\text{cls}}$ acts as the hidden feature for classifying $\mathbf{x}$. For simpler notation, we denote a function $\text{Vit}:X\rightarrow Y$ to summarize all computations involved in the ViT including the last softmax function, where $\mathbf{x}\in X$ and $y\in Y$. The parameters of the ViT are estimated using maximum likelihood estimation (MLE), or equivalently, through minimizing the cross entropy (CEloss) between the prediction and the ground truth:
@@
III-C Test-time Predictions
In this subsection, the probabilistic model proposed in Sec. III-A is used to predict the class given an image $\mathbf{x}$ sampled from $p_{\text{test}}(\mathbf{x})$. Once trained, a 3-phase procedure is proposed to obtain the final prediction as illustrated in Fig. 2 (a).
@@ III-C Test-time Predictions
Given the estimated expected response vector $\mathbf{y}\in\mathbb{R}^{A}$ obtained by averaging all $K\times M$ samples from the mixture model $p(\mathbf{y}|\mathbf{x},\Theta)$, let $\mathbf{y}^{a}$ represent its value in the $a$-th dimension. The probability of the final prediction being the class indexed by $a$ is then calculated in the softmax form of the Brier score. Here we follow the formula and the hyperparameter $\iota\in\operatorname{\mathbb{R}}$ introduced in diffusion-based classifiers by Han et al. [40]:
@@
IV Experiments and Results
Figure 4: Illustration of the Tuberculosis chest X-ray dataset under different perturbations.
We evaluate the proposed method on two medical imaging benchmarking datasets: Tuberculosis chest X-ray dataset [22] and the ISIC Melanoma skin cancer dataset [23]. The Tuberculosis chest X-ray dataset consists of X-ray images of 3500 patients with Tuberculosis and 3500 patients without Tuberculosis. The ISIC Melanoma skin cancer dataset contains lesion images of 5105 patients with malignant skin cancer and 5500 patients with benign tumor. A range of baseline methods are chosen for comparisons and these cover a variety of different architectures:
@@ IV Experiments and Results
  • Comparison with Existing Ensemble Methods:  To provide a comprehensive evaluation, the proposed ensemble method is compared against state-of-the-art ensemble deep learning methods specifically designed for robust medical image classification, including deep tree training of convolutional ensembles (DTT) [18], improved convolutional ensembles (ICNN-Ensemble) [46] and dynamic-weighted ensembles (DWE) [19].
Experiment Configuration Details.  For the chest X-ray dataset, the split for the image-label pairs in training/validation/testing set is $5670/630/700$. For the ISIC skin cancer dataset, the split for the image-label pairs in training/validation/testing set is $7605/1000/2000$. In both datasets, all images have binary labels. The test sets are balanced. Several empirical choices were made: 5 mixture components ($K=5$), and each were sampled 20 times ($M=20$). For the diffusion model, 1000 time-steps ($T=1000$) were chosen, with a noise schedule of $\beta_{1}=10^{-4}$, $\beta_{T}=0.02$. For the probability simplex mapping, $\iota=0.1737$ was set for the X-ray dataset, and $\iota=0.3162$ for the skin cancer dataset. The transformer encoder blocks in LaDiNE are extracted from ViT-B [37], and all mapping networks are implemented by multilayer perceptrons (MLPs) with 3 hidden layers. All baseline models and LaDiNE were trained from scratch until loss convergence. For the baseline methods that require selecting ensemble members, the procedures in the original papers were followed exactly as described. The implementation of LaDiNE, including the code for models as well as training/evaluation scripts, is made publicly available.
IV-A Results without Perturbations
When no perturbations are performed on the input images, performance is very good for all methods. For the chest X-ray dataset, when presented with clean inputs (i.e., without simulated covariate shifts), all methods achieve classification accuracies that exceed $99.00\%$. For the skin cancer dataset, the proposed method achieves accuracies of $94.18\%$ on clean inputs, on par with the best-performing method ResNet-18 which attains accuracies of $95.02\%$.
IV-B Results Under Perturbations
@@
TABLE I: Comparison in classification accuracy (%) with state-of-the-art methods on two benchmark datasets with unseen input perturbations. Methods are categorised into four groups, from top to bottom: (i) Models based on convolutional neural networks (CNNs). (ii) Models based on Transformers; hybrid models with CNNs and Transformers (ConViT-B, MedViT-B), and diffusion models (CARD). (iii) Ensemble learning methods.
@@
Results.  The experimental results for the robustness experiments on both datasets can be found in Table. I. The results presented indicate the means and standard deviations of the classification accuracies (in percentages) over three runs. The overall trend indicates that LaDiNE consistently outperforms other methods across almost all perturbations, highlighting its effectiveness and robustness in handling noisy and perturbed images. In particular, LaDiNE shows superior performance on both datasets under high noise levels ($\mathbb{g}=1.00$), and demonstrates the highest robustness among all tested models overall. Traditional models such as ResNet-18 and ResNet-50 exhibit poor performance under Gaussian noise, with accuracies dropping to $50\%$  [^2] Note that ResNet-50 consistently provides predictions of “Healthy” for all input images perturbed with Gaussian noise ($\mathbb{g}=1.00$)., indicating a failure to generalize under noisy conditions.
When handling lower-resolution images, LaDiNE achieves the highest accuracies, particularly in the chest X-ray dataset ($98.90\%$), outperforming other methods such as ViT-B and ConViT-B. Although model performance typically declines with lower resolution input images, the transformer-based models (i.e. the second group of models in Table. I), such as ViT-B and DeiT-B, maintain relatively high accuracies as compared to CNN-based models. When handling lower-contrast input images, LaDiNE and ResNet-50 lead in robustness on the ISIC dataset, with LaDiNE achieving $93.14\%$ accuracy. Transformer-based models generally show robust performance across varying contrast levels, with SEViT and ConViT-B also performing well.
IV-C Results Under Adversarial Attacks
@@
Results.  The results presented in Table. II demonstrate the performance of various methods under adversarial attacks using FGSM, PGD, and AutoPGD algorithms. LaDiNE, consistently outperforms other models across both datasets and all attack types, indicating superior robustness to adversarial attacks.
@@
TABLE II: Comparison in classification accuracy (%) with state-of-the-art methods on two benchmarks with adversarial attacks.
@@
IV-D Results on the Quality of Prediction Confidence
@@
IV-E Quantifying Instance-level Uncertainty
Quantifying the reliability of a model’s predictions is critical in clinical settings, where the consequences of presenting incorrect predictions can be significant. In order to maintain trust in the system, the model should quantify the level of uncertainty in each prediction, with the goal of being correct when confident, and uncertain when incorrect [61]. In this fashion, clinicians can focus their review on cases where the model is less certain, thereby improving decision-making and fostering trust in the system. This sections shows results in the quantification of LaDiNE’s prediction uncertainties given an image instance under covariate shifts.
Recall that LaDiNE provides $K\times M$ prediction vectors $\hat{\mathbf{y}}$ for a given image $\mathbf{x}$ (see Sec. III-C). These prediction vectors are denoted as a set $\mathcal{S}$ and a set of scalars are defined: $\mathcal{S}_{a}:=\{\hat{\mathbf{y}}_{i}^{a}\mid\hat{\mathbf{y}}_{i}\in\mathcal{S}\}$ (note that $\hat{\mathbf{y}}_{i}^{a}$ denotes the value of $\hat{\mathbf{y}}_{i}$ in the $a$-th dimension). Uncertainties are measured with respect to the consistency among the samples predictions. To this end, two methods are used:
@@ IV-E Quantifying Instance-level Uncertainty
TABLE III: Results of uncertainty measures under covariate shift (Images with Gaussian noise injection $\mathbb{g}=1.00$).
@@
In order to examine the power of the method in challenging contexts, the input images are perturbed significantly with Gaussian noise ($\mathbb{g}=1.00$), see example images in Fig. 6. This level of Gaussian noise results in it being challenging to differentiate healthy images from tuberculosis images. Results shown in Table. III illustrate the instance-level uncertainty estimates for the LaDiNE’s predictions on different classes, specifically focusing on the distinction between correct and incorrect predictions. The first thing to note is that the results reflect how the challenging context results in high uncertainties for the healthy class predictions. LaDiNE appropriately expresses the uncertainty in these challenging cases, informing clinicians to review those uncertain instances more carefully. On the other hand, LaDiNE is more certain when correctly predicting tuberculosis, which demonstrates the model’s robustness in detecting true unhealthy cases.
TABLE IV: Comparison in classification accuracy (Acc. %) and expected calibration error (ECE) with different designs. $M$ indicates the number of sampling times from the distribution. We investigate the effectiveness of inferred latent variable and diffusion model in improving classification accuracy and confidence calibration.
@@ IV-F Ablation Studies
As shown in Table. IV, the complete version of LaDiNE (design 4) achieves the highest accuracy for all testing sets, the lowest relative accuracy drop under noisy conditions, and the lowest Expected Calibration Error (ECE), providing support and justification for the design elements chosen.
@@
Results on Selection of Mixture Components.  Another key design choice in our method is the selection of mixture components, in other words, the TE hierarchy. Table. V shows the performance of the proposed method with $K=3,4,\dotsc,7$ (we draw 20 samples per ensemble member). When $K=5$, the classification accuracy is the highest for both datasets. When $K$ is smaller, there is insufficient discrimination in the extracted features which therefore results in low classification accuracy. On the other hand, as $K$ increases, the inner structure of feature representations from the different hierarchies becomes too complex and may therefore result in a performance drop.
TABLE V: Comparison in classification accuracy (%) with different choice of mixture components.
@@
Examining the results in Table. V and the findings in Fig. 1, one can observe a trade-off between informativeness and invariance in selecting early TE blocks’ representation. Specifically:
@@
V Limitations and Future Works
One of the limitations of our approach is the increased computational cost associated with diffusion models, which require iterative denoising processes. Although the denoising process is on $\operatorname{\mathbb{R}}^{A}$ instead of the whole image space ($A$ is the number of classes), it is still limited to tasks that do not require time-sensitive responses. To address this, future work could explore faster sampling techniques, such as Denoising Diffusion Implicit Models (DDIM) [63] or consistency models [64], which can reduce the number of required iterations and computational overhead while maintaining comparable performance. In addition to incorporating accelerated sampling techniques, the computation of ensemble members can be parallelized to further reduce the latency and improve efficiency. By distributing the ensemble’s workload across multiple processing units (e.g., multiple graphical processing units), inference times can be significantly reduced, making the approach more efficient and practical for various applications.

### v5

@@ 3.2 Training
Step 1 Training TE Blocks.  The parameters of the TE blocks are learned during the training of the ViT. As discussed previously, the input $\mathbf{x}$ is initially transformed into an embedding and subsequently processed through $L$ Transformer Encoder (TE) blocks. The class token generated by the final TE block is used for classifying $\mathbf{x}$. For simpler notation, we denote a function $\text{Vit}:X\rightarrow Y$ to summarize all computations involved in the ViT including the last softmax function, where $\mathbf{x}\in X$ and $y\in Y$. The parameters of the ViT are estimated using maximum likelihood estimation (MLE), or equivalently, through minimizing the cross entropy (CEloss) between the prediction and the ground truth:
@@
3.3 Test-time Predictions
In this subsection, the probabilistic model proposed in Sec. 3.1 is used to predict the class given an image $\mathbf{x}$ sampled from $p_{\text{test}}(\mathbf{x})$. Once trained, a 3-phase procedure is proposed to obtain the final prediction as illustrated in Fig. 2 (a).
@@ 3.3 Test-time Predictions
Given the estimated expected response vector $\mathbf{y}\in\mathbb{R}^{A}$ obtained by averaging all $K\times M$ samples from the mixture model $p(\mathbf{y}|\mathbf{x},\Theta)$, let $\mathbf{y}^{a}$ represent its value in the $a$-th dimension. The probability of the final prediction being the class $a$ is then calculated in the weighted-softmax form of the Brier score [47]. Specifically, we follow the mapping function introduced in diffusion-based classifiers by Han et al. [40], where the hyperparameter $\iota$ controls the sharpness of the probability distribution:
@@
4 Experiments and Results
Figure 4: Illustration of the Tuberculosis chest X-ray dataset under different perturbations (noisy $\mathbb{g}=0.10$, low resolution $w=8.00$, contrast $r=0.70$, FGSM adversarial attack $\varepsilon=0.03$).
We evaluate the proposed method on two medical imaging benchmarking datasets: Tuberculosis chest X-ray dataset [23] and the ISIC Melanoma skin cancer dataset [24]. The Tuberculosis chest X-ray dataset consists of X-ray images of 3500 patients with Tuberculosis and 3500 patients without Tuberculosis. The ISIC Melanoma skin cancer dataset contains lesion images of 5085 patients with malignant skin cancer and 5480 patients with benign tumor. A range of baseline methods are chosen for comparisons and these cover a variety of different architectures:
@@ 4 Experiments and Results
  • Comparison with Existing Ensemble Methods:  To provide a comprehensive evaluation, the proposed ensemble method is compared against state-of-the-art ensemble deep learning methods specifically designed for medical image classification, including deep tree training of convolutional ensembles (DTT) [19], improved convolutional ensembles (ICNN-Ensemble) [50] and dynamic-weighted ensembles (DWE) [20].
Experiment Configuration Details.  For the chest X-ray dataset, the split for the image-label pairs in training/validation/testing set is $5670/630/700$. For the ISIC skin cancer dataset, the split for the image-label pairs in training/validation/testing set is $7605/1000/1960$. In both datasets, all images have binary labels. The test sets are balanced. Several empirical choices were made: 5 mixture components ($K=5$), and each were sampled 20 times ($M=20$). For the diffusion model, 1000 time-steps ($T=1000$) were chosen, with a noise schedule of $\beta_{1}=10^{-4}$, $\beta_{T}=0.02$. For the probability simplex mapping, we determine the optimal value of $\iota$ by minimizing the ECE on the validation set, using the Nelder-Mead method [51], specifically, $\iota=0.1737$ was set for the X-ray dataset, and $\iota=0.3162$ for the skin cancer dataset. The transformer encoder blocks in LaDiNE are extracted from ViT-B [42], and all mapping networks are implemented by multilayer perceptrons (MLPs) with 3 hidden layers. All CDMs and $\text{Enc}(\cdot)$ are implemented with MLPs with 3 hidden layers, the conditioning in the CDM (time and image $\mathbf{x}$ conditioning) is implemented with element-wise product. All baseline models and LaDiNE were trained from scratch until loss convergence. For the baseline methods that require selecting ensemble members, the procedures in the original papers were followed exactly as described. The implementation of LaDiNE, including the code for models as well as training/evaluation scripts, is made publicly available [^2] The code repository: https://github.com/xingbpshen/nested-diffusion.
4.1 Results without Perturbations
When no perturbations are performed on the input images, performance is very good for all methods. For the chest X-ray dataset, when presented with clean inputs (i.e., without simulated covariate shifts), all methods achieve classification accuracies that exceed $99.00\%$. For the skin cancer dataset, the proposed method achieves accuracies of $94.18\%$ on clean inputs, on par with the best-performing method ResNet-18 which attains accuracies of $95.02\%$ [^3] For the clean chest X-ray dataset, ViT-B achieves $99.75\%$ accuracy, LaDiNE achieves $99.90\%$ accuracy. For the clean skin cancer dataset, ViT-B achieves $92.93\%$ accuracy..
4.2 Results Under Perturbations
@@
Table 1: Comparison in classification accuracy (%) with state-of-the-art methods on two benchmark datasets with unseen input perturbations. Methods are categorized into four groups, from top to bottom: (i) Models based on convolutional neural networks (CNNs). (ii) Models based on Transformers. (iiI) hybrid models with CNNs and Transformers (ConViT-B, MedViT-B), and diffusion models (CARD). (iv) Ensemble learning methods.
@@
Results.  The experimental results for the robustness experiments on both datasets can be found in Table. 1. The results presented indicate the means and standard deviations of the classification accuracies (in percentages) over three runs. The overall trend indicates that LaDiNE consistently outperforms other methods across almost all perturbations, highlighting its effectiveness and robustness in handling noisy and perturbed images. In particular, LaDiNE shows superior performance on both datasets under high noise levels ($\mathbb{g}=1.00$), and demonstrates the highest robustness among all tested models overall. Traditional models such as ResNet-18 and ResNet-50 exhibit poor performance under Gaussian noise, with accuracies dropping to $50\%$  [^4] Note that ResNet-50 consistently provides predictions of “Healthy” for all input images perturbed with Gaussian noise ($\mathbb{g}=1.00$)., indicating a failure to generalize under noisy conditions.
When handling lower-resolution images, LaDiNE achieves the highest accuracies, particularly in the chest X-ray dataset ($98.90\%$). Although model performance typically declines with lower resolution input images, the transformer-based models (i.e. the second group of models in Table. 1), such as ViT-B and DeiT-B, maintain relatively high accuracies as compared to CNN-based models. When handling lower-contrast input images, LaDiNE and ResNet-50 lead in robustness on the ISIC dataset, with LaDiNE achieving $93.14\%$ accuracy. Transformer-based models generally show robust performance across varying contrast levels, with SEViT and ConViT-B also performing well.
4.3 Results Under Adversarial Attacks
@@
Results.  The results presented in Table. 2 demonstrate the performance of various methods under adversarial attacks using FGSM, PGD, and AutoPGD algorithms. LaDiNE, consistently outperforms other models across both datasets and all attack types, indicating superior robustness to adversarial attacks.
@@
Table 2: Comparison in classification accuracy (%) with state-of-the-art methods on two benchmarks with adversarial attacks.
@@
4.4 Results on the Quality of Prediction Confidence
@@
4.5 Quantifying Instance-level Uncertainty
Quantifying the reliability of a model’s predictions is critical in clinical settings, where the consequences of presenting incorrect predictions can be significant. In order to maintain trust in the system, the model should quantify the level of uncertainty in each prediction, with the goal of being correct when confident, and uncertain when incorrect [66]. In this fashion, clinicians can focus their review on cases where the model is less certain, thereby improving decision-making and fostering trust in the system. This section shows results in the quantification of LaDiNE’s prediction uncertainties given an image instance under covariate shifts.
Recall that LaDiNE provides $K\times M$ prediction vectors $\hat{\mathbf{y}}$ for a given image $\mathbf{x}$ (see Sec. 3.3). These prediction vectors are denoted as a set $\mathcal{S}$ and a set of scalars are defined: $\mathcal{S}_{a}:=\{\hat{\mathbf{y}}_{i}^{a}\mid\hat{\mathbf{y}}_{i}\in\mathcal{S}\}$ (note that $\hat{\mathbf{y}}_{i}^{a}$ denotes the value of $\hat{\mathbf{y}}_{i}$ in the $a$-th dimension). Uncertainties are measured with respect to the consistency among the samples predictions. To this end, two methods are used:
@@ 4.5 Quantifying Instance-level Uncertainty
Table 3: Results of uncertainty measures under covariate shift (Images with Gaussian noise injection $\mathbb{g}=1.00$).
@@
In order to examine the power of the method in challenging contexts, the input images are perturbed significantly with Gaussian noise ($\mathbb{g}=1.00$), see example images in Fig. 6. This level of Gaussian noise results in it being challenging to differentiate healthy images from tuberculosis images. Results shown in Table. 3 illustrate the instance-level uncertainty estimates for the LaDiNE’s predictions on different classes, specifically focusing on the distinction between correct and incorrect predictions. The first thing to note is that the results reflect how the challenging context results in high uncertainties for the healthy class predictions. LaDiNE appropriately expresses the uncertainty in these challenging cases, informing clinicians to review those uncertain instances more carefully. On the other hand, LaDiNE is more certain when correctly predicting tuberculosis, which demonstrates the model’s robustness in detecting true unhealthy cases.
Table 4: Comparison in classification accuracy (Acc. %) and expected calibration error (ECE) with different designs. $M$ indicates the number of sampling times from the distribution. We investigate the effectiveness of inferred latent variable and diffusion model in improving classification accuracy and confidence calibration.
@@ 4.6 Ablation Studies
As shown in Table. 4, the complete version of LaDiNE (design 4) achieves the highest accuracy for all testing sets, the lowest relative accuracy drop under noisy conditions, and the lowest Expected Calibration Error (ECE), providing support and justification for the design elements chosen.
@@
Results on Selection of Mixture Components.  Another key design choice in our method is the selection of mixture components, in other words, the TE hierarchy. Table. 5 shows the performance of the proposed method with $K=3,4,\dotsc,7$ (we draw 20 samples per ensemble member). When $K=5$, the classification accuracy is the highest for both datasets. When $K$ is smaller, there is insufficient discrimination in the extracted features which therefore results in low classification accuracy. On the other hand, as $K$ increases, the inner structure of feature representations from the different hierarchies becomes too complex and may therefore result in a performance drop.
Table 5: Comparison in classification accuracy (%) with different choice of mixture components.
@@
Examining the results in Table. 5 and the findings in Fig. 1, one can observe a trade-off between informativeness and invariance in selecting early TE blocks’ representation. Specifically:
@@
5 Limitations and Future Works
One of the limitations of our approach is the increased computational cost associated with diffusion models, which require iterative denoising processes. Although diffusion models are expensive, the denoising process is on $\mathbb{R}^{A}$ instead of on the whole image space ($A$ is the number of classes). Inference time is still reasonable for many medical imaging tasks. In our experiments, the inference time (on one NVIDIA A100 GPU) per image is approximately $98.23$ ms for LaDiNE (we sequentially sample $K\times M$ times, where $K=5$ and $M=20$), compared to $0.54$ ms for the ViT-B baseline. Future work could explore faster sampling techniques, such as Denoising Diffusion Implicit Models (DDIM) [68] or consistency models [69], which can reduce the number of required iterations and computational overhead while maintaining comparable performance. In addition to incorporating accelerated sampling techniques, the computation of ensemble members (implemented by $K$ mapping networks and denoising networks) can be parallelized to further reduce the latency and improve efficiency. By distributing the ensemble’s workload across multiple processing units (e.g., multiple graphical processing units), inference times can be significantly reduced, making the approach more efficient and practical for various applications.